A car is traveling with a velocity v on a circular road of radius r. If it is decreasing it`s speed at the rate of a, then the magnitude of resultant acceleration will be.......

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
 
 

The instantaneous acceleration ar pointing towards the center of the circle is called radial acceleration or centripetal acceleration.                      The radial component causes a change in direction and its magnitude is given by ar = v2/r, where r is the radius of curvature at the point in question.

 

The tangential component causes a change in speed and its magnitude is given by at = dv/dt = a m/sec2 (given)

 

The tangential acceleration of the car has a constant magnitude of at = a m/sec2

 

The components aand at are perpendicular to each other and the magnitude of a is given by

 

resultant accln. a2 = ar2 + at2

 

                         resultant accln. (ares)         =[ a2+(v2/r)]1/2

                                                                       =[ a2 +(v4/r2) ]1/2

 

Regards

Arun(askIITians forum expert)